# TopoLM: brain-like spatio-functional organization in a topographic language model

### v1

@@ 5.1 Clusters selective for verbs and nouns
Using fMRI data from Hauptman et al. (2024) (Appendix B), we find verb- and noun-selective clusters in the left hemisphere (Figure 3A and Figure 6), thus replicating their results. To quantify the ‘degree’ of clustering in these maps, we use Moran’s $I$ with Queen contiguity (i.e. $\ell^{\infty}$ radius, following the distance metric used to train TopoLM), a common measure of spatial autocorrelation, ranging from $-1$ to $1$ (see Appendix A for details). The group level effects map yields $I=0.96$ ($p<0.001$), indicating strong clustering.
Figure 3: Brain-like verb- and noun-selective clusters in TopoLM. (a) fMRI data from Hauptman et al. (2024) points to verb- and noun-selective regions in the human cortex with strong clustering (Moran’s $I=0.96$). (b) Quantification of clustering. Relative to high clustering in the brain (green dashed line), the non-topographic baseline shows limited clustering (orange). The topographic model shows moderate clustering at the unit level (light blue) and strong clustering when simulating fMRI sampling (dark blue). On stimuli from Moseley & Pulvermüller (2014), another neuroimaging study on verb/noun-selectivity (fMRI data not available) we find qualitatively similar results. (c) Exemplary model maps (last MLP layer) showing the verb-/noun contrast (red-blue) in response to stimuli from (Hauptman et al., 2024). The non-topographic baseline shows no clustering while the topographic model develops verb- and noun-selective clusters.
@@
To investigate whether similar verb- and noun-selective clusters emerge in TopoLM, we extracted the model activations in response to the same stimuli as Elli et al. (2019) and Hauptman et al. (2024). Contrasting model activations to verb and noun stimuli yields clear verb- and noun-selective clusters across layers in our topographic model (verb: red / noun: blue, Figure 3A, Figure 9). Similar to our analysis of the neuroimaging data, we quantified this observation using Moran’s $I$. The non-topographic baseline model yields a low degree of clustering ($I=0.11$). Contrast maps in TopoLM show a strong degree of clustering ($I=0.48$) which is further increased when applying fMRI-like readout sampling ($I=0.81$). Applying the same sampling to the non-topographic model also increases clustering, but it remains substantially less brain-like than the topographic model ($I=0.60$, Figure 3B,C).
@@ 5.2 Clusters selective for concrete, but not abstract verb-noun contrasts
Clustering in TopoLM. We presented the original experimental stimuli to TopoLM and computed contrasts between abstract and concrete verbs and nouns (see Appendix C for stimulus details). Rather than comparing response profiles in anatomically defined subregions as in Moseley & Pulvermüller (2014) (since anatomical regions are not well-defined in TopoLM), we instead explore whether lexical class-selective clustering emerges between concrete and abstract words. Consistent with brain data, we find clustering of verb- and noun-selective model units when concrete words were used, but find only very weak or no clustering for the same contrast using abstract words. We quantified this impression using again Moran’s $I$: concrete words yield substantially higher verb-/noun-clustering than abstract words ($I=0.8$ vs. $I=0.23$ in unthresholded maps for TopoLM, $t$-test: $p<0.001$, Figure 4, light blue). We obtain qualitatively similar results when simulating the fMRI sampling process before computing contrasts (Figure 4, dark blue). Additionally, we observed overall low degrees of clustering in non-topographic baseline models and no evidence for a difference in the degree of clustering when concrete vs. abstract words were used for computing the verb-/noun-contrast (Figure 4, orange).
@@ 6 Downstream Performance and Brain Alignment
Some topographic vision models sacrifice task performance for the sake of spatial organization, which leads to diminished utility as a model of brain function (Lee et al., 2020). As a control, we thus evaluate TopoLM’s performance on linguistic knowledge (BLiMP), downstream task (GLUE), and brain alignment benchmarks (Brain-Score), and compare it to the non-topographic baseline model.
@@
BLiMP (Warstadt et al., 2020) is a multi-task linguistic benchmark consisting of minimal pairs (acceptable / unacceptable) that probe specific knowledge of linguistic phenomena; we measure overall accuracy as the proportion of pairs to which the model assigns higher probability to the grammatically acceptable sentence. Models are evaluated without fine-tuning. Brain-Score Language (Schrimpf et al., 2021b; 2020) is a set of benchmarks that measure how well models align with the human language system. Following prior work, we train a ridge regression model to predict brain activity from model representations, using the same stimuli as in neuroimaging studies. Representations are taken from each transformer block where spatial loss is applied, including both attention and MLP layers. For both models, we select the output layer with the best alignment across datasets. The final score is the Pearson correlation between actual and predicted brain activity, averaged over 10 cross-validation folds and over four brain-recording datasets (Blank et al., 2014; Fedorenko et al., 2016; Pereira et al., 2018; Tuckute et al., 2024, Appendix D). GLUE (Wang et al., 2018) is a multi-task benchmark for downstream performance on tasks like entailment and sentiment analysis. Following common practice, we fine-tune models for each task. During fine-tuning, we continue to use the same weighted combination of task loss (cross-entropy or MCC) and spatial correlation loss with $\alpha=2.5$. We set batch size 64, dropout 0.1, and learning rate $2\times 10^{-5}$; all other hyperparameters are the same as pretraining. We use early stopping after 3 successive increases in validation loss, keeping the checkpoint before this increase.
@@
Table 1: Results from evaluation of TopoLM and the non-topographic baseline model (‘non-topo’) on BLiMP, GLUE, and Brain-Score. The non-topographic model outperforms TopoLM on BLiMP (by 5 points) and Brain-Score (by 2 points) out-of-the-box, but TopoLM outperforms the non-topographic control on GLUE (by 3 points) after fine-tuning.
@@ Results.
We find that introducing spatial correlation loss does not strongly affect task performance or brain alignment (Figure 1). On BLiMP, we find slight decreases in performance between TopoLM and the non-topographic baseline across all subtasks, with a 5 point average decrease overall. On Brain-Score, we find a 2 point average decrease overall, with TopoLM notably outperforming the non-topographic baseline on some subtasks (see Appendix D). On GLUE, across most subtasks we see an increase in performance from TopoLM compared to the baseline model, with a 3 point average increase overall. We hypothesize this increase might be due the spatial loss term serving as an additional regularization against overfitting.
@@ 7 Discussion
Here, we present a new topographic Transformer language model based on a spatial smoothness constraint, and showed that it predicts key patterns of spatio-functional organization from the neuroimaging literature. In particular, we demonstrated that TopoLM has a spatially organized core language system where it exhibits brain-like uniformity of response profiles across clusters. TopoLM also exhibits brain-like noun-verb selective clusters, which are specific to concrete over abstract words. This additional spatial correspondence to the human brain seems to come at virtually no cost to performance and functional brain alignment.
@@
TopoLM is a feed-forward model of linguistic processing with multiple layers. Since the topographic maps are specific to each layer, there is as such no coherent tissue across the entire system as is the case for the brain. Transitions and space between clusters is thus likely not well accounted for. This also limits eventual modeling of ‘BMI’ tissue perturbations, such as micro-stimulation, since the spread of current ends at one topographic map (out of several).
While we evaluated our model on three distinct human brain datasets related to topographic linguistic processing, these are far from capturing the topography of language in the brain. New neuroimaging experiments are needed to this effect, which might further validate or also falsify the model—we see this as an opportunity for future experiments.
@@
Taken together, our results suggest that the spatial smoothness principle leads to topographic organization that is consistent with the spatio-functional organization of linguistic processing in the brain.
@@ 8 Acknowledgments
We thank Miriam Hauptman for providing fMRI data and for useful discussion. We also thank Greta Tuckute and members of the EPFL NeuroAI lab for helpful feedback. Neil Rathi was supported by a Summer@EPFL Fellowship.
@@ Appendix A Moran’s $I$ and Spatial Autocorrelation
Both Elli et al. (2019) and Hauptman et al. (2024) are based on fMRI recordings in response to the same task and stimuli—similarity judgments of auditorily presented words—and both investigate the relationship between sensory input and conceptual representations in the human cortex. Elli et al. (2019) ($n=13$) provide evidence suggesting that lexico-semantic information about verbs and nouns is represented in partially non-overlapping cortical networks (Figure 1A in (Elli et al., 2019)).
Hauptman et al. (2024) presented the same auditory stimuli to 21 congenitally blind and 22 sighted subjects and investigated how neural representations of sensory concepts develop independent of sensory experience; here, we only include data from the 22 sighted subjects in our analysis. Similarly to Elli et al. (2019), Hauptman et al. (2024) find distinct verb- and noun-selective regions (e.g. Figure 3A in Hauptman et al. (2024); Figure 6 below). For scanning parameters, and details on the fMRI preprocessing pipeline, see Hauptman et al. (2024). Here, we investigate whether similar non-overlapping cortical verb- and noun-selective regions emerge in TopoLM.
@@
In each 2-second trial, subjects assessed the similarity of two auditorily presented words from the same verb or noun subcategory (e.g., bird or mammal for nouns, light or sound for verbs). Four trials, all using stimuli from the same subcategory, form a mini-block, followed by a 10-second baseline condition. This yielded a total of $288$ pairs of nouns and $288$ pairs of verbs. [^3] Each pair was presented with verbs in the infinitive (with ‘to’) and nouns in the definite (with ‘the’). Removing these determiners when feeding stimuli through the model results in visually more distinct selectivity clusters; see Figure 9.
To evaluate each vertex’s selectivity for a mini-block, we ran a general linear model with one predictor per mini-block. Contrasting verb and noun mini-blocks produced a $t$-value for each vertex, indicating selectivity toward verbs vs. nouns. As in Hauptman et al. (2024), we based our analysis on data from the left hemisphere (Frost, 1999; Szaflarski et al., 2006), which is more involved in language processing; this is standard in work on language neuroscience.
@@
In each trial, 18 participants from Moseley & Pulvermüller (2014) silently read a single word from one of four categories: abstract nouns (e.g. clue, guide), concrete nouns (e.g. rice, goose), abstract verbs (e.g. heal, dwell), and concrete verbs (e.g. knit, poke). In total, there were $40$ stimuli in each category for a total of $160$ words. For additional information on the experimental design and neuroimaging parameters, please see Moseley & Pulvermüller (2014)).
@@ Appendix D Brain-Score Language Datasets and Performance
Figure 5: Brain-Score brain alignment performance. We tested TopoLM and the non-topographic control on Brain-Score language. TopoLM outperforms the control in Blank2014 and Tuckete2024, but performs slightly worse at Pereira2018 and Fedorenko2016 ($t$-test performed for each benchmark separately: $p<0.05$). For each benchmark, we sampled from 10 cross-validation loops to compute the bootstrapped 95% confidence intervals (black bars). When results are averaged across the 4 benchmarks (last panel ‘Brain-Score’), we don’t find evidence for a significant difference between TopoLM and its control ($t$-test: $p>0.05$).
@@
This dataset captures ECoG signals from five participants as they read sentences consisting of eight words, presented one word at a time for either 450 or 700 ms. Following Schrimpf et al. (2021b), we focus on 52 out of 80 sentences presented to all participants.
@@ Appendix E Supplementary Figures
Figure 8: Consistent response profiles across language-selective clusters in TopoLM. Using stimuli from Fedorenko et al. (2024b), we extract response profiles to 4 linguistic categories across language-selective clusters. Importantly, the observed response profiles are consistent across language-selective clusters. For brevity, here, we only show clusters with more than 100 units.

### v2

@@ 5.1 Clusters selective for verbs and nouns
Using fMRI data from Hauptman et al. (2024) (Appendix B), we find verb- and noun-selective clusters in the left hemisphere (Figure 3A and Appendix Figure 7), thus replicating their results. To quantify the ‘degree’ of clustering in these maps, we use Moran’s $I$ with Queen contiguity (i.e. $\ell^{\infty}$ radius, following the distance metric used to train TopoLM), a common measure of spatial autocorrelation, ranging from $-1$ to $1$ (see Appendix A for details). The group level effects indicate strong clustering ($I=0.96$, $p<0.001$).
@@
To investigate whether similar verb- and noun-selective clusters emerge in TopoLM, we extracted the model activations in response to the same stimuli as Elli et al. (2019) and Hauptman et al. (2024). We find that verb and noun-selective clustering emerges across layers in TopoLM after contrasting activations to verb and noun stimuli (verb: red / noun: blue, Figure 3A, Figure 10). Similar to our analysis of the neuroimaging data, we quantified this observation using Moran’s $I$. The non-topographic baseline model yields a low degree of clustering ($I=0.11$). Contrast maps in TopoLM show a strong degree of clustering ($I=0.48$) which is further increased when applying fMRI-like readout sampling ($I=0.81$, Figure 3B). Applying the same sampling to the non-topographic model also increases clustering, but it remains substantially less brain-like than the topographic model ($I=0.60$, Appendix Figure 10).
Clustering in Topoformer-BERT.
Applying the same procedure to Topoformer-BERT, we find no evidence for noun-verb selective clustering, with few units coming out significant in the noun-verb contrast ($10.61\%$ of units; see Figure 13). However, we do find that, before thresholding for significance ($p<0.05$), the model does exhibit a high degree of clustering competitive with TopoLM (Moran’s $I=0.66$ before sampling, $0.85$ after sampling). In other words, though local connectivity constraint induces clustering in the model, these clusters do not match the spatio-functional organization of the brain. This impression is confirmed by additional anaylsis using a variant of Moran’s I that only considers units with significant t-values (Figure 14).
@@ 5.2 Clusters selective for concrete, but not abstract verb-noun contrasts
Clustering in TopoLM. We presented the original experimental stimuli to TopoLM and computed contrasts between abstract and concrete verbs and nouns (see Appendix C for stimulus details). Rather than comparing response profiles in anatomically defined subregions as in Moseley & Pulvermüller (2014) (since the model lacks defined ‘anatomical’ regions), we explore whether lexical class-selective clustering emerges between concrete and abstract words. Consistent with brain data, we find clustering of verb- and noun-selective model units in concrete stimuli, but find only very weak or no clustering for the same contrast in abstract stimuli. We quantified this impression using again Moran’s $I$: concrete words yield substantially higher verb-/noun-clustering than abstract words ($I=0.8$ vs. $I=0.23$ in unthresholded maps for TopoLM, Figure 4, light blue). We obtain qualitatively similar results when simulating the fMRI sampling process before computing contrasts (Figure 4, dark blue). Additionally, we observed overall low degrees of clustering in non-topographic baseline models and no evidence for a difference in the degree of clustering when concrete vs. abstract words were used for computing the verb-/noun-contrast (Figure 4, orange).
Clustering in Topoformer-BERT.
We again apply the same procedure to Topoformer-BERT and find no evidence for noun-verb selective clustering, with no units coming out significant in the noun-verb contrast (see Figure 13). Again, before thresholding, we find high clustering (concrete-concrete: Moran’s $I=0.60$ before sampling, $I=0.84$ after; abstract-abstract: $I=0.61$ before sampling, $I=0.84$ after); yet since none of these units come out significant, we fail to find evidence for brain-like spatio-functional organization in Topoformer-BERT (for details, see Figure 14).
@@ 6 Downstream Performance and Brain Alignment
Some topographic vision models sacrifice task performance for the sake of spatial organization, leading to diminished utility as a model of brain function (Lee et al., 2020). As a control, we evaluate TopoLM’s performance on linguistic knowledge (BLiMP), downstream task (GLUE), and brain alignment benchmarks (Brain-Score), compared to the non-topographic baseline model.
@@
BLiMP (Warstadt et al., 2020) is a multi-task linguistic benchmark consisting of minimal pairs (acceptable / unacceptable) that probe knowledge of linguistic phenomena; Models are evaluated without fine-tuning. Brain-Score Language (Schrimpf et al., 2021; 2020) is a set of benchmarks that measure how well models align with the human language system. We train a ridge regression model to predict brain activity from model representations, using the same stimuli as in the neuroimaging studies. Representations are taken from each transformer block with spatial loss (attention and MLP). For both models, we select the output layer with the best alignment across datasets. The final score is the Pearson correlation between actual and predicted brain activity, averaged over 10 cross-validation folds and over four brain-recording datasets (Blank et al., 2014; Fedorenko et al., 2016; Pereira et al., 2018; Tuckute et al., 2024, Appendix D). GLUE (Wang et al., 2018) is a multi-task benchmark for downstream performance on tasks like entailment and sentiment analysis. We fine-tune models for each task. During fine-tuning, we use the same weighted combination of task loss (cross-entropy or MCC) and spatial correlation loss with $\alpha=2.5$. We set batch size 64, dropout 0.1, and learning rate $2\times 10^{-5}$; all other hyperparameters are the same as during pretraining. We use early stopping after 3 successive increases in validation loss, keeping the checkpoint before this increase.
@@
Table 1: Results from evaluation of TopoLM and a non-topographic baseline (‘non-topo’) on BLiMP, GLUE, and Brain-Score. The baseline model outperforms TopoLM on BLiMP (5 pts) and Brain-Score (2 pts), but TopoLM outperforms the non-topographic control on GLUE (3 pts).
@@ Results.
We find that introducing spatial correlation loss does not strongly affect task performance or brain alignment (Table 1). On BLiMP, we find slight decreases in performance between TopoLM and the non-topographic baseline across all subtasks, with a 5 point average decrease overall. On Brain-Score, we find a 2 point average decrease overall, with TopoLM outperforming the non-topographic baseline on some subtasks (see Appendix D, Figure 6). On GLUE, across most subtasks we see an increase in performance from TopoLM compared to the baseline model, with a 3 point average increase overall. We hypothesize this increase might be due the spatial loss term serving as additional regularization against overfitting.
@@ 7 Discussion
Here, we presented a new topographic Transformer language model based on a spatial smoothness constraint and showed that it predicts key patterns of spatio-functional organization from the neuroimaging literature. We demonstrated that TopoLM has a spatially organized core language system exhibiting brain-like uniformity of response profiles across clusters. TopoLM also shows brain-like noun-verb selective clusters, specific to concrete over abstract words. This spatial correspondence to the human brain comes at virtually no cost to performance or functional brain alignment.
@@
Comparison with Topoformer-BERT.
We also observe evidence for spatial organization with noun-verb contrasts in TopoLM, but not BinHuraib et al. (2024)’s Topoformer-BERT. However, Fedorenko et al. (2010)’s stimuli do reveal a spatially organized language network in Topoformer-BERT. This asymmetry is likely in part because the model induces topography only in the attention layers, which primarily track relationships between tokens: while Fedorenko et al. (2010)’s stimuli consists of entire sentences, the noun-verb contrast stimuli are one or two word items.
However, it’s important to note that—while this comparison is a useful baseline—our primary goal is to ask whether the TDANN spatial smoothness principle generalizes to language, not to build a topographic language model that “improves” on certain benchmarks. Topoformer asks a similar question using a different guiding principle (local connectivity), but these models do not necessarily “compete” with one another: indeed, TopoLM’s spatial smoothness and Topoformer’s explicit local connectivity are both closely related to the principle of wiring length minimization.
@@
TopoLM is a feed-forward model of linguistic processing with multiple layers. Since the topographic maps are specific to each layer, there is as such no coherent tissue across the entire system as in the the brain. This also limits eventual modeling of ‘BMI’ tissue perturbations, such as micro-stimulation, since the spread of current ends at one topographic map (out of several).
@@
Taken together, our results suggest that the spatial smoothness principle leads to topographic organization consistent with the spatio-functional organization of linguistic processing in the brain.
@@ 8 Acknowledgments
We thank Miriam Hauptman and Marina Bedny for providing fMRI data and for useful discussion. We also thank Greta Tuckute, members of the EPFL NeuroAI lab, and members of the Stanford CLiMB Lab for helpful feedback. Neil Rathi was supported by a Summer@EPFL Fellowship.
@@ Appendix A Moran’s $I$ and Spatial Autocorrelation
Both Elli et al. (2019) and Hauptman et al. (2024) are based on fMRI recordings in response to the same task and stimuli—similarity judgments of auditorily presented words—and both investigate the relationship between sensory input and conceptual representations in the human cortex. Elli et al. (2019) ($n=13$) provide evidence suggesting that lexico-semantic information about verbs and nouns is represented in partially non-overlapping cortical networks (Figure 1A in Elli et al. (2019)). The data from Hauptman et al. (2024) are available at OPENICPSR.
Hauptman et al. (2024) presented the same auditory stimuli to 21 congenitally blind and 22 sighted subjects and investigated how neural representations of sensory concepts develop independent of sensory experience; here, we only include data from the 22 sighted subjects in our analysis. Similarly to Elli et al. (2019), Hauptman et al. (2024) find distinct verb- and noun-selective regions (e.g. Figure 3A in Hauptman et al. (2024); Figure 7 below). For scanning parameters, and details on the fMRI preprocessing pipeline, see Hauptman et al. (2024).
@@
In each 2-second trial, subjects assessed the similarity of two auditorily presented words from the same verb or noun subcategory (e.g., bird or mammal for nouns, light or sound for verbs). Four trials, all using stimuli from the same subcategory, form a mini-block, followed by a 10-second baseline condition. This yielded a total of $288$ pairs of nouns and $288$ pairs of verbs [^9] Each pair was presented with verbs in the infinitive (with ‘to’) and nouns in the definite (with ‘the’). Removing these determiners when feeding stimuli through the model results in visually more distinct selectivity clusters; see Figure 10..
To evaluate each vertex’s selectivity for a mini-block, we ran a general linear model with one predictor per mini-block. Contrasting verb and noun mini-blocks produced a $t$-value for each vertex, indicating selectivity toward verbs vs. nouns. As in Hauptman et al. (2024), we based our analysis on data from the left hemisphere, which is more involved in language processing (Frost, 1999; Szaflarski et al., 2006); this is standard in work on language neuroscience.
@@
In each trial, 18 participants from Moseley & Pulvermüller (2014) silently read a single word from one of four categories: abstract nouns (e.g. clue, guide), concrete nouns (e.g. rice, goose), abstract verbs (e.g. heal, dwell), and concrete verbs (e.g. knit, poke). In total, there were $40$ stimuli in each category for a total of $160$ words. For additional information on the experimental design and neuroimaging parameters, please see Moseley & Pulvermüller (2014).
@@ Appendix D Brain-Score Language Datasets and Performance
Figure 5: Brain-Score brain alignment performance (linear predictivity). We tested TopoLM and the non-topographic control on Brain-Score language using linear predictivity to estimate alignment. TopoLM outperforms the control in Blank2014 and Tuckete2024, but performs slightly worse at Pereira2018 and Fedorenko2016 ($t$-test performed for each benchmark separately: $p<0.05$). For each benchmark, we sampled from 10 cross-validation loops to compute the bootstrapped 95% confidence intervals (black bars). When results are averaged across the 4 benchmarks (last panel ‘Brain-Score’), we don’t find evidence for a significant difference between TopoLM and its control ($t$-test: $p>0.05$).
Figure 6: Brain-Score brain alignment performance (RSA). We tested TopoLM and the non-topographic control on Brain-Score language using representational similarity analysis (RSA) to estimate alignment. TopoLM outperforms the control in Pereira2018, but performs worse at Blank2014 and Fedorenko2016 ($t$-test performed for each benchmark separately: $p<0.05$). For each benchmark, we sampled from 10 cross-validation loops to compute the bootstrapped 95% confidence intervals (black bars). When results are averaged across the 3 benchmarks (last panel ‘Brain-Score’), we find evidence for a significant difference between TopoLM and its control ($t$-test: $p>0.05$). However, this result needs to be interpreted with caution as it is dominated by the results from Fedorenko2016 which only show very large absolute values, because the underlying data used for normalization is very noisy, i.e. the noise ceiling is very low.
@@
This dataset captures ECoG signals from five participants as they read sentences consisting of eight words, presented one word at a time for either 450 or 700 ms. Following Schrimpf et al. (2021), we focus on 52 out of 80 sentences presented to all participants.
@@ Appendix E Supplementary Figures
Figure 9: Consistent response profiles across language-selective clusters in TopoLM. Using stimuli from Fedorenko et al. (2024), we extract response profiles to 4 linguistic categories across language-selective clusters. Importantly, the observed response profiles are consistent across language-selective clusters. For brevity, here, we only show clusters with more than 100 units.
@@
Figure 12: Verb-noun clusters in a TopoLM without randomly permuted positions. As described in Footnote 2, we trained an equivalent model without first randomly permuting the spatial encoding of each layer. We evaluate this model on the stimuli from Hauptman et al. (2024) and Moseley & Pulvermüller (2014). We see that the spatial activation pattern propagates through the entire network, as the model no longer has to learn unique spatial maps for each layer in order to minimize spatial loss.
Figure 13: Topoformer-BERT Evaluations. Contrast maps (with fMRI-like readout sampling) for Topoformer-BERT (BinHuraib et al., 2024) on stimuli from Fedorenko et al. (2010), Hauptman et al. (2024), and Moseley & Pulvermüller (2014). We plot contrasted activations after multiplication by the key matrix and after the entire attention layer, as this is where local connectivity is applied. We note that few units come out significant for the stimuli from Hauptman et al. (2024) and Moseley & Pulvermüller (2014); however, we identify some clusters in Fedorenko et al. (2010)’s language localization task.
Figure 14: TopoLM comparison of clustering with baseline model. We compared TopoLM to an additional baseline model (Topoformer BERT, BinHuraib et al. (2024)) with regard to clustering. Most t-values across contrasts in multiple datasets (Hauptman et al. (2024); Moseley & Pulvermüller (2014)) are not significant in Topoformer BERT (after applying FDR correction as has been done in TopoLM). Thus, we estimated autocorrelation using a variant of Moran’s I only considering ”islands” of units with significant t-values for a given contrast (and averaging across islands while not considering island size to obtain a single estimate per map). We computed ”Islands Moran’s I” for both TopoLM and Topoformer BERT. A) Using this metric, we find strong clustering in TopoLM (Island’s Moran’s I = 0.52 without simulated fMRI sampling, and Island’s Moran’s I = 0.83 with sampling), numbers matching our results when computing Standard Moran’s I (I = 0.53 without simulated fMRI sampling, I = 0.85 with sampling). In contrast, we observe patterns of dispersion or neither clear dispersion, nor autocorrelation in Topoformer BERT (Island’s Moran’s I = -0.51 without simulated fMRI sampling, and Island’s Moran’s I = -0.14 with sampling). B-E) With the exception of one contrast (see panel E), TopoLM results based on data from Moseley & Pulvermüller (2014) using Islands Moran’s I qualitatively match the results obtained by using Standard Moran’s I. In contrast, in Topoformer BERT autocorrelation using Islands Moran’s I cannot be computed for the contrasts in Moseley & Pulvermüller (2014) as t-values underlying the contrast maps are not significant. F) Legend.